Show that the light from a source which is moving away from us with a speed of

0.6c has twice the wavelength it has in the frame of reference of the source.



1
Expert's answer
2021-06-13T11:29:24-0400

We can use the relativistic Doppler Effect equation in this problem.

For a source of light moving away from the observer, the wavelength of light measured by the observer will be:

"λ=λ_0 \sqrt{ \frac{1+β}{1-β} }"

λ0 is the wavelength of light in frame of reference of the source

We can compute the β value as

"β = \frac{v}{c} \\

β = \frac{0.6c}{c} \\

β =0.6"

Substitute the β value:

"λ=λ_0 \sqrt{ \frac{1+0.6}{1-0.6} } \\

λ=λ_0 \sqrt{4} \\

λ=2λ_0"

The light from a source which is moving away from us with a speed of 0.6c has twice the wavelength it has in the frame of reference of the source.